# Stable Nonconvex-Nonconcave Training via Linear Interpolation

### 12 Jan 2024 v2

@@ Abstract
This paper presents a theoretical analysis of linear interpolation as a principled method for stabilizing (large-scale) neural network training. We argue that instabilities in the optimization process are often caused by the nonmonotonicity of the loss landscape and show how linear interpolation can help by leveraging the theory of nonexpansive operators. We construct a new optimization scheme called relaxed approximate proximal point (RAPP), which is the first 1-SCLI method to achieve last iterate convergence rates for $\rho$-comonotone problems while only requiring $\rho>-\tfrac{1}{2L}$. The construction extends to constrained and regularized settings. By replacing the inner optimizer in RAPP we rediscover the family of Lookahead algorithms for which we establish convergence in cohypomonotone problems even when the base optimizer is taken to be gradient descent ascent. The range of cohypomonotone problems in which Lookahead converges is further expanded by exploiting that Lookahead inherits the properties of the base optimizer. We corroborate the results with experiments on generative adversarial networks which demonstrates the benefits of the linear interpolation present in both RAPP and Lookahead.
@@ 1 Introduction
Table 1: Overview of last iterate results with our contribution highlighted in blue. Prior to this work there existed no rates for 1-SCLI schemes handling $\rho$-comonotone problems with $\rho\in(-\nicefrac{{1}}{{2L}},\infty)$ and no global convergence guarantees for Lookahead beyond bilinear games.
@@
This work focuses on 1-SCLI schemes (Arjevani et al., 2015; Golowich et al., 2020), whose update rule only depends on the previous iterate in a time-invariant fashion. Another approach to establishing last iterate is Halpern-type methods with an explicit scheme developed in Lee & Kim (2021) for cohypomonotone problems and later extended to the constrained case in Cai et al. (2022) (c.f. LABEL:app:relatework).
@@ 4 Inexact Krasnosel’skiĭ-Mann iterations
An operator $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ is said to be quasi-nonexpansive if
@@ 6 Last iterate under cohypomonotonicity
1:$z^{0}\in\mathbb{R}^{n}$ $\lambda\in(0,1)$, $\gamma\in(\lfloor-2\rho\rfloor_{+},\nicefrac{{1}}{{L}})$
@@ 10 Acknowledgements
Thanks to Leello Dadi and Stratis Skoulakis for helpful discussion. This work was supported by Google. This work was supported by Hasler Foundation Program: Hasler Responsible AI (project number 21043). This work was supported by the Swiss National Science Foundation (SNSF) under grant number 200021_205011. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n° 725594 - time-data).

### 14 Mar 2024 v4

@@ Abstract
This paper presents a theoretical analysis of linear interpolation as a principled method for stabilizing (large-scale) neural network training. We argue that instabilities in the optimization process are often caused by the nonmonotonicity of the loss landscape and show how linear interpolation can help by leveraging the theory of nonexpansive operators. We construct a new optimization scheme called relaxed approximate proximal point (RAPP), which is the first explicit method without anchoring to achieve last iterate convergence rates for $\rho$-comonotone problems while only requiring $\rho>-\tfrac{1}{2L}$. [^1] The range of RAPP can be extended to $\rho>-\tfrac{1}{L}$ by using the weaker notion of conic nonexpansiveness as first exploited in Alacaoglu et al. (2024) to derive first-order complexity results (cf. LABEL:app:conic). The construction extends to constrained and regularized settings. By replacing the inner optimizer in RAPP we rediscover the family of Lookahead algorithms for which we establish convergence in cohypomonotone problems even when the base optimizer is taken to be gradient descent ascent. The range of cohypomonotone problems in which Lookahead converges is further expanded by exploiting that Lookahead inherits the properties of the base optimizer. We corroborate the results with experiments on generative adversarial networks which demonstrates the benefits of the linear interpolation present in both RAPP and Lookahead.
@@ 1 Introduction
Table 1: Overview of last iterate results with our contribution highlighted in blue. Prior to this work there existed no rates for explicit schemes without anchoring handling $\rho$-comonotone problems with $\rho\in(-\nicefrac{{1}}{{2L}},\infty)$ and no global convergence guarantees for Lookahead beyond bilinear games.
@@
This work focuses on 1-SCLI schemes (Arjevani et al., 2015; Golowich et al., 2020), whose update rule only depends on the previous iterate in a time-invariant fashion. Another approach to establishing last iterate is Halpern-type methods with an explicit scheme developed in Lee & Kim (2021) for cohypomonotone problems and later extended to the constrained case in Cai et al. (2022) (cf. LABEL:app:relatework).
@@ 4 Inexact Krasnosel’skiĭ-Mann iterations
An operator $T:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is said to be quasi-nonexpansive if
@@ 6 Last iterate under cohypomonotonicity
1:$z^{0}\in\mathbb{R}^{d}$ $\lambda\in(0,1)$, $\gamma\in(\lfloor-2\rho\rfloor_{+},\nicefrac{{1}}{{L}})$
@@ 10 Acknowledgements
Thanks to Ahmet Alacaoglu, Weiqiang Zheng, Leello Dadi and Stratis Skoulakis for helpful discussion. This work was supported by Google. This work was supported by Hasler Foundation Program: Hasler Responsible AI (project number 21043). This work was supported by the Swiss National Science Foundation (SNSF) under grant number 200021_205011. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n° 725594 - time-data).